IONS AT THE SURFACE OF A ROOM-TEMPERATURE IONIC LIQUID

\underline{CHERRY S. SANTOS}, SELIMAR RIVERA-RUBERO, SERGEY DIBROV, and STEVEN BALDELLI, Department of Chemistry, University of Houston, Houston, TX 77204.

The orientation and interfacial location of both the cation and anion of the ionic liquids, 1-butyl-3-methylimidazolium methylsulfate, [BMIM][MS] and 1-butyl-3-methylimidazolium methylsulfonate, [BMIM][CH_{3}SO_{3}], have been determined by sum-frequency generation vibrational spectroscopy and X-ray crystallography. The results indicate that both the cation and anion occupy the first layer at the gas-liquid interface. Further, the methyl groups of the cation and anion are directed away from the liquid phase into the gas phase.